the width of a rectangle is 12 ft less than the length. the area of the rectangle is 540 square ft. Find the dimensions of the rectangle
Found 2 solutions by oberobic, mananth:
Answer by oberobic(2304)

L = W+12
A = L*W = 540 sq ft
A = (W+12)*W = 540
W^2+12W-540 = 0
(W +30)(W-18) = 0
W = -30 or 18
Negative width is nonsense, so
W = 18
L = 18+12 = 30
Answer: Dimensions are length 30, width 18.
Done.

Answer by mananth(16949)

length = x
width = (x-12)
Area = L*W
Area = x(x-12)=540
x^2-12x=540
x^2-12x-540=0
x^2-30x+18x-540=0
x(x-30)+18(x-30)=0
(x-30)(x+18)=0
x=30 which is positive
length = 30 ft
width = 18 feet
m.ananth@hotmail.ca